Y=-3x + 8
Y=-5x - 2

Answer by ankor@dixie-net.com(22740)

Y=-3x + 8
Y=-5x - 2
y = y, therefore
-3x + 8 = -5x - 2
-3x + 5x = -2 - 8
2x = -10
x = -5
Find y using the 1st equation, check solutions using the 2nd equation